A number consists of two digits.The digit at the ten's place is two times the digit at the unit's place.The number formed by reversing the digits is 27 less than the original number.Find the original number.
Answer by mananth(12270)


let x be in the tens place and y in the units place.
x=2y
10y+x=10x+y -27
9y-9x=-27
plug the value of x in the equation
9y-18y=-27
-9y=-27
y=3
x=6
the number is 63